If a rectangle has a length that is twice its width, what is the length of the rectangle if the area is 162 ft.
Answer by satyareddy22(84)

Let l=length,b=width
l=2b
area=l*b=162
2b*b=162
=>2b^2=162
=>b^2=81=>b=9
=>l=2b=2*9=18
l=18,b=9